OpenStudy (anonymous):

Find x A. 9 B. 18 C. 9√3 D. 3√3

OpenStudy (♪chibiterasu):

The equation?

OpenStudy (♪chibiterasu):

We will be using Tangent. x is the hypotenuse and 9 is the opposite. 30 is the theta: \[tan30 = \frac {9}{x}\]

OpenStudy (♪chibiterasu):

\[tan30 \approx 0.577\]

OpenStudy (♪chibiterasu):

\[0.577 \approx \frac {9}{x}\] Multiply x on both sides. then divide tan30 on both sides and solve. \[(0.577)(x) \approx 9\] \[x \approx \frac {9}{0.577}\] \[x \approx 15.59\]

OpenStudy (♪chibiterasu):

Hmm, I think I did something wrong.

OpenStudy (anonymous):

haha thanks for helping me tho. I am having trouble with this unit

OpenStudy (anonymous):

do you have any other idea on how to do this?

OpenStudy (♪chibiterasu):

Yeah I'm kind of confused as well. >.<

OpenStudy (anonymous):

haha ok thanks anyways. Do you know of anyone else who can help?

OpenStudy (anonymous):

@thomaster

OpenStudy (♪chibiterasu):

OOOOH I see what I did wrong. I needed to use sine and not tangent.

OpenStudy (♪chibiterasu):

sin30 = 0.5 9/0.5 = 18

OpenStudy (♪chibiterasu):

Ha, silly me. :P

OpenStudy (anonymous):

haha thank you so much!

OpenStudy (anonymous):

can you help with one more?

OpenStudy (♪chibiterasu):

sure

OpenStudy (♪chibiterasu):

Using SOH CAH TOA, do you know which triginomic ratio we need to use? Cosine, sine or tangent? 

OpenStudy (anonymous):

wow this is making me see how truly bad i am at this... im not really sure... promise im not just tryin to get answers haha

OpenStudy (anonymous):

oh and the options are A.9

OpenStudy (anonymous):

B. 9√2 C.4.5√2 D. 18√2

OpenStudy (♪chibiterasu):

SOH means: sine = opossite/hypotenuse CAH means: cosine = adjacent/hypotenuse TOA means: tangent = opposite/adjacent They are abbreviations. We have the opposite measurement and are looking for the adjacent. So which ratio do we need to use?

OpenStudy (anonymous):

TOA?

OpenStudy (♪chibiterasu):

Yes. We are going to use the tangent. So: \[tan45 = \frac {18}{x}\] Do you know how to find x?

OpenStudy (anonymous):

ugh. no not quite

OpenStudy (♪chibiterasu):

Well we learned in previous math lessons that to find a missing variable you need to isolate the unknown variable. So in this case we need to find a way to move the "x" so that it is by itself. Are you understanding so far?

OpenStudy (anonymous):

ya i understand what you r saying

OpenStudy (♪chibiterasu):

But we need to know what tan45 is. But we can't do that calculation so quickly, which is why we have calculators. Do you know how to find tan45 with a calculator? :3

OpenStudy (anonymous):

no

OpenStudy (♪chibiterasu):

Some calculators are different. Do you have one at home?

OpenStudy (anonymous):

ya

OpenStudy (♪chibiterasu):

Ok, are there tan, cos and sin buttons on it?

OpenStudy (anonymous):

is it 1? i put 45 tan

OpenStudy (♪chibiterasu):

Yep!

OpenStudy (anonymous):

yay i actually got something right! haha

OpenStudy (♪chibiterasu):

So now we have: \[1 = \frac {9}{x}\] So now let's isolate the x. What's the first step?

OpenStudy (anonymous):

multiplying 9 and 1?

OpenStudy (♪chibiterasu):

Nope, because x is the denominator then we must multiply x on both sides. That gives us 1x on the left, but what will happen on the right?

OpenStudy (anonymous):

oh ok..... 9x? i have no idea

OpenStudy (♪chibiterasu):

Since it's already being divided by x, it will be canceled out by the multiplication. So now we have: \[1x = 9\] What is the last step?

OpenStudy (anonymous):

oh ok i see.

OpenStudy (anonymous):

umm devide 1 and 9 so its 9

OpenStudy (♪chibiterasu):

Yes. So x = 9. And to check, we have to make sure that tan45 = 9/9 tan45 = 1 x = 9 1 = 9/9 1 = 1 :D

OpenStudy (anonymous):

yay thank you so much!!!

OpenStudy (♪chibiterasu):

Aaaah I mixed up 18 with 9 somehow. .-.

OpenStudy (♪chibiterasu):

Hold on xD

OpenStudy (♪chibiterasu):

1 = 18/x 1x = 18 x = 18/1 x = 18 1 = 18/18 1 = 1 Yeah xD

OpenStudy (anonymous):

so i was right with 9?

OpenStudy (♪chibiterasu):

Omg I'm so stupid 18 was the hypotenuse .-.